OpenStudy (anonymous):

A new computer store offers a special early-bird prize give-away for its grand opening. The first 12 customers to enter the store between 7 a.m. and 8 a.m. will receive gift certificates worth various dollar amounts. If 18 customers gather to enter the store between those hours, how many ways are there to arrange the twelve prize winners?

OpenStudy (anonymous):

@Brostep0s

OpenStudy (anonymous):

n = 18, r = 12. Read the language: 18 customers have gathered, so definitely the first 12 customers will be from these 18. 12 are to be chosen from these 18.

OpenStudy (anonymous):

So, how would I write the answer exactly?

OpenStudy (anonymous):

18P12=18/6=18*17*16*15*14*13=

OpenStudy (anonymous):

@BradSuga did you get what you were looking for?

OpenStudy (anonymous):

Yes I did thank you!